OpenStudy (anonymous):

PLEASE HELP I WILL FAN AND METAL PLEASE! given ABC with A(-4,-2) B(4,4) and C(18,-8) write and equation of the line containing the altitude that passes through B in standard form.

OpenStudy (anonymous):

You could use A,C to write the equation of a line, then check B is on it.

OpenStudy (anonymous):

can you help me work it out because i don't understand this at all...

OpenStudy (anonymous):

@redcan ^

OpenStudy (anonymous):

Which form of the formula of a line do you want to start with?

OpenStudy (anonymous):

point slope form? \[y-y_1 = m(x-x_1)\]

OpenStudy (anonymous):

um sure

OpenStudy (anonymous):

Ok, how do you find m, the slope?

OpenStudy (anonymous):

y2-y1/x2-x1 ?

OpenStudy (anonymous):

yep, so in our case m = (-2 - 8)/(-4-18) = ?

OpenStudy (anonymous):

-3/11?

OpenStudy (anonymous):

\[y−y_1=(-3/11)(x−x_1)\] now sub in y1,x1 from point A of C

OpenStudy (anonymous):

*or